Two hallways, of widths 'a' and 'b', meet at right angles. What is the greatest possible length of a ladder which can be carried horizontally around the corner ?


1
Expert's answer
2021-03-09T06:58:37-0500

Solution:

Let's draw a figure according to the question and label it.



Let AB be the ladder with length L. C is the point of intersection of hallways.

So, "L=AC+BC"

Given, "BE=b, AD=a"

Let "CD=y, CE=x"

Then, both triangles ACD and CEB are right-angled triangles.

By pythagoras theorem, in triangle ACD,

"AD^2+CD^2=AC^2
\\\Rightarrow a^2+y^2=AC^2"

"\Rightarrow AC=\sqrt {a^2+y^2}"

By pythagoras theorem, in triangle CEB,

"CE^2+EB^2=BC^2
\\\Rightarrow x^2+b^2=BC^2"

"\Rightarrow BC=\sqrt {x^2+b^2}"

Next, "\angle CAD=\angle BCE" [Corresponding angles]

We take their tangent as: "\dfrac{CD}{AB}=\dfrac{BE}{CE}"

"\Rightarrow \dfrac{y}{a}=\dfrac{b}{x}"

"\Rightarrow y=\dfrac{ab}{x}" ...(i)

Now, "L=AC+BC=\sqrt {a^2+y^2}+\sqrt {x^2+b^2}"

"=\sqrt {a^2+(\dfrac{ab}{x})^2}+b\sqrt {(\dfrac{x}{b})^2+1}" [Using (i)]

Put "\dfrac{x}{b}=t"

"L=a\sqrt {1+(\dfrac{1}{t})^2}+b\sqrt {(t)^2+1}"

"=a\sqrt {\dfrac{t^2+1}{t^2}}+b\sqrt {t^2+1}"

"=\dfrac at\sqrt {t^2+1}+b\sqrt {t^2+1}"

"=\sqrt {t^2+1}(\dfrac at+b)" ...(ii)

Now differentiating w.r.t. "t" ,

"L'=(\sqrt {t^2+1})'(\dfrac at+b)+\sqrt {t^2+1}(\dfrac at+b)'"

"=(\dfrac{1}{2\sqrt {t^2+1}}\times 2t)(\dfrac at+b)+\sqrt {t^2+1}(\dfrac {-a}{t^2}+0)"

"=(\dfrac{t}{\sqrt {t^2+1}})(\dfrac at+b)-(\dfrac {a\sqrt {t^2+1}}{t^2})"

Put "L'=0"

"\Rightarrow (\dfrac{t}{\sqrt {t^2+1}})(\dfrac at+b)-(\dfrac {a\sqrt {t^2+1}}{t^2})=0"

"\Rightarrow (\dfrac{t}{\sqrt {t^2+1}})(\dfrac at+b)=\dfrac {a\sqrt {t^2+1}}{t^2}"

"\Rightarrow {a+bt}=\dfrac {a(t^2+1)}{t^2}"

"\Rightarrow t^2( {a+bt})= a(t^2+1)"

"\Rightarrow at^2+bt^3= a(t^2+1)
\\ \Rightarrow bt^3=a
\\ \Rightarrow t^3=\dfrac ab \Rightarrow t=(\dfrac ab)^{1/3}"

Put this value of t in (ii)

"L=\sqrt {[(\dfrac ab)^{1/3}]^2+1}\times(\dfrac a{(\dfrac ab)^{1/3}}+b)"

"=\sqrt {[(\dfrac ab)^{2/3}]+1}\times(\dfrac a{(\dfrac {a^{1/3}}{b^{1/3}})}+b)"

"=\sqrt {\dfrac {a^{2/3}+b^{2/3}}{b^{2/3}}}\times( {a^{2/3}b^{1/3}}+b)"

"= {\dfrac {\sqrt{a^{2/3}+b^{2/3}}}{b^{1/3}}}\times b^{1/3}( {a^{2/3}}+b^{2/3})"

"=( {a^{2\/3}}+b^{2\/3})^{3\/2}"

This is required length of the ladder.